A box with a surface area of 94 inches squared is 3 inches high and 5 inches long, how wide is the box?? how do i figure this out... please help
Answer by checkley77(12844)

94=2(3*5)+2(3*X)+2(5*X)
94=2*15+6X+10X
94=30+16X
16X=94-30
16X=64
X=64/16
X=4 ANS. FOR THE WIDTH.
PROOF:
94=2*15+2(3*4)+2(5+4)
94=30+2*12+2*20
94=30+24+40
94=94